# Exact model reduction for conditional quantum dynamics

### v1

@@ 2 Compute the output algebra $\mathscr{A}={\rm alg}(\mathscr{N}^{\perp})$;
3 Compute the factorizations of $\mathbb{E}|_{\mathscr{A}}$, $\mathcal{R}$ and $\mathcal{J}$;
@@
Let us consider a CE ($(\{\mathcal{M}_{k}\},\mathcal{C})$ with observation algebra $\mathscr{A}={\rm alg}(\mathscr{N}^{\perp})$. Let $\mathbb{E}|_{\mathscr{A}}$ be the conditional expectation onto $\mathscr{A}$ and $\mathcal{R}$ and $\mathcal{J}$ its CPTP factorizations. Then, fixing $\Phi=\mathcal{R}$, the reduced CE defined over $\widecheck{\mathscr{A}}:={\rm Im}(\mathcal{R})$, of conditional dynamics ˇ${\widecheck{\mathcal{M}}}_{k}:=\mathcal{R}\mathcal{M}_{k}\mathcal{J}$, output map $\widecheck{\mathcal{C}}:=\mathcal{C}\mathcal{J}$ and initial condition $\tau(0)=\mathcal{R}[\rho_{0}]$ solves Problem 1, that is for all sequences $k_{0:t}$ and for all $t\geq 0$ and $\rho_{0}\in\mathfrak{D}(\mathcal{H})$
@@
for all $\rho_{0}\in\mathfrak{D}(\mathcal{H})$ and sequences $k_{0:t}$ with $t\geq 0$. By noticing that $\mathscr{N}^{\perp}\subseteq\mathscr{A}$ and using Propositon 2 we can conclude the proof. ∎
Using the same arguments as in [16], one can prove that the closure to an algebra of $\mathscr{N}^{\perp}$, orthogonal to $\mathscr{N}$ with respect to the Hilbert-Schmidt inner product ,$_{HS}$, provides the smallest algebra containing $\mathscr{N}^{\perp}$ that admits a CPTP projection, even when one allows for alternative definitons of operator product and orthogonality notions.
@@ IV-C Reduction of measurements and dynamics, separately
In the previous sections, we presented how to reduce a CE to a reduced-order one, obtaining a set of conditional dynamics that are capable of reproducing the output of interest conditioned on the outcome of the measurement. In many cases of interest, the conditional dynamic emerges as the composition of an unconditional CPTP evolution and a generalized measurement [30]. Under suitable assumtpions, it is possible to obtain reduced models for both the CPTP dynamical map and the measurements, separately. In the following we thus assume that the conditional dynamics $\mathcal{M}_{k}$ are composed of a CPTP evolution map $\mathcal{E}$ following the conditioning induced by the generalized measuerement $\mathcal{K}_{k}(\cdot):=M_{k}\cdot M_{k}^{\dagger}$ with $\sum_{k\in\Omega}M_{k}^{\dagger}M_{k}=\mathbb{1}$, i.e. $\mathcal{M}_{k}(\cdot)=\mathcal{E}\circ\mathcal{K}_{k}$. This ordering choice has been made to keep consistency with previous works [13, 15] but similar results can be derived if the evolution precedes the measurement.
@@
Notice that Assumption (A1) implies Assumption (A2): by definition $\mathscr{N}$ is $\mathcal{M}_{k}$-invariant for all $k$, hence for $X\in\mathscr{N}$, we have $\mathcal{E}[X]=\sum_{k}\lambda_{k}\mathcal{M}_{k}[X]\in\mathscr{N}$. This generalizes the results presented in [13, equation 3.3] and [15, Lemma 4] to the case of stochastic output matrices (not only deterministic ones). Assumption (A2) however, does not imply (A1).
@@
Under any of the assumptions (A1)-(A4), the reduced CE ({ˇ$(\{\widecheck{\mathcal{M}}_{k}\},\widecheck{\mathcal{C}})$ with conditional dynamics ˇ$\widecheck{\mathcal{M}}_{k}=\widecheck{\mathcal{E}}\widecheck{\mathcal{K}}_{k}$ where $\widecheck{\mathcal{K}}_{k}=\mathcal{R}\mathcal{K}_{k}\mathcal{J}$ and ˇ$\widecheck{\mathcal{E}}=\mathcal{R}\mathcal{E}\mathcal{J}$, $\widecheck{\mathcal{C}}=\mathcal{C}\mathcal{J}$ and $\Phi=\mathcal{R}$ solves Probelm 1.
@@ V-A Measured quantum walks
In [15, Proposition 3], we proved that, for commutative algebras, including the effects of conditioning in the definition of $\mathscr{N}^{\perp}$, can only enlarge it, implying that the reduced model that includes the effects of conditioning is always at least as large as the unconditional reduced model. This is no longer true for non-commutative algebras, i.e. quantum systems, where it is possible that the conditional reduction leads to a smaller model than the unconditional one. The following is a pedagogical yet fundamental example of this fact.
@@
It is worth remarking that this type of situation, i.e. conditioning induces smaller reduced models, is not possible in the classical case. In fact [13] proves that the conditional observable space always contains the unconditional one.
@@ V-B Measured Ising spin chain
Let us consider a quantum system composed of $N\geq 4$ qubits disposed on a line [^1] We denote the operator living in $\mathfrak{B}(\mathcal{H})$ and acting as the Pauli matrix $\sigma_{j}$ with $j\in\{0,x,y,z\}$ on the $k$-th spin, with $k=1,\dots,N$ and as the identity on the rest is denoted with $\sigma_{j}^{(k)}$, i.e. $\sigma_{j}^{(k)}:=\otimes_{i=1}^{k-1}\sigma_{0}\otimes\sigma_{j}\otimes_{j=1}^{N-k}\sigma_{0}\in\mathfrak{B}(\mathcal{H})$., i.e. $\mathcal{H}\simeq(\mathbb{C}^{2})^{\otimes N}$. Let us also assume to be able to perform projective measurement of the observable $\sigma_{z}$ on the last spin, i.e. $\sigma_{z}^{(N)}$ and that, at each step, this measurement is performed with probability $1-p$ with $p\in[0,1]$ and it is left untouched with probability $p$. Under these assumptions we can model the measurement with three super operators $\mathcal{K}_{-1}(\cdot)=p\mathbb{1}\cdot\mathbb{1}$, $\mathcal{K}_{0}(\cdot)=(1-p)\Pi_{0}\cdot\Pi_{0}$ with $\Pi_{0}:=\mathbb{1}_{2^{N-1}}\otimes\left|0\middle>\middle<0\right|$ and $\mathcal{K}_{1}(\cdot)=(1-p)\Pi_{1}\cdot\Pi_{1}$ with $\Pi_{1}=\mathbb{1}_{2^{N-1}}\otimes\left|1\middle>\middle<1\right|$. Between two measurements the system evolves with a unitary evolution $\mathcal{E}(\cdot)=U\cdot U^{\dagger}$ with $U=e^{-iH}$ where $H$ is an Ising Hamiltonian, i.e.
@@
with coupling strength $\delta$.
For this example, we are interested in reproducing the reduced state on the first qubit, i.e. $\tau(t)={\rm tr}_{\overline{0}}(\rho(t))$ for all initial conditions $\rho_{0}\in\mathfrak{D}(\mathcal{H})$, which can be reconstructed from the expectation of $\left\{\sigma_{q}^{(1)}\right\}_{q=0,x,y,z}$, representing our observables of interest. We next derive the reduced model in two cases of interest: $p=0$ and $10<p<$ 1 and we denote the spaces and maps of interest with the index 00 and $p$ to distinguish them.
@@
with $V_{k}:\mathcal{H}\to\mathcal{H}_{k}$ isometries of the correct dimensions and where $\mathcal{H}_{k}=\mathcal{H}_{S,k}\otimes\mathcal{H}_{F,k}\simeq\mathbb{C}^{4}\otimes\mathbb{C}^{2^{N-3}}$. The reduced evolution map takes the form ˇ$\widecheck{\mathcal{M}}_{p}(\cdot)=\widecheck{U}_{p}\cdot\widecheck{U}_{p}^{\dagger}$ where, $\widecheck{U}_{p}=\exp\left(-i\delta[\sigma_{0}\otimes\sigma_{0}\otimes\sigma_{z}+\sigma_{z}\otimes\sigma_{x}\otimes\sigma_{0}]\right)$ and the reduced state on the first qubit is retrieved by $\tau(t)=\sum_{k=0}^{1}{\rm tr}_{\overline{0}}[\xi_{k}(t)]$. Moreover, conditioning for the reduced model is associated to $\widecheck{\mathcal{K}_{0}}(\cdot)=p\mathbb{1}_{8}\cdot\mathbb{1}_{8}$, $\widecheck{\mathcal{K}_{0}}(\cdot)=(1-p)\hat{\Pi}_{0}\cdot\widecheck{\Pi}_{0}$ and $\widecheck{\mathcal{K}_{1}}(\cdot)=(1-p)\widecheck{\Pi}_{1}\cdot\widecheck{\Pi}_{1}$ with $\widecheck{\Pi}_{0}=\frac{\mathbb{1}_{8}+\mathbb{1}_{4}\otimes\sigma_{x}}{2}$ and $\widecheck{\Pi}_{1}=\frac{\mathbb{1}_{8}-\mathbb{1}_{4}\otimes\sigma_{x}}{2}$.

### v2

@@ 2 Compute the output algebra $\mathscr{A}={\rm alg}(\mathscr{N}^{\perp})$;
3 Compute the factorization of $\mathbb{E}|_{\mathscr{A}}$, $\mathcal{R}$ and $\mathcal{J}$;
@@
Let us consider a CE ($(\{\mathcal{M}_{k}\},\mathcal{C})$ with observation algebra $\mathscr{A}={\rm alg}(\mathscr{N}^{\perp})$. Let $\mathbb{E}|_{\mathscr{A}}$ be the conditional expectation onto $\mathscr{A}$ and $\mathcal{R}$ and $\mathcal{J}$ its CPTP factorization. Then, fixing $\Phi=\mathcal{R}$, the reduced CE defined over $\widecheck{\mathscr{A}}:={\rm Im}(\mathcal{R})$, of conditional dynamics ˇ${\widecheck{\mathcal{M}}}_{k}:=\mathcal{R}\mathcal{M}_{k}\mathcal{J}$, output map $\widecheck{\mathcal{C}}:=\mathcal{C}\mathcal{J}$ and initial condition $\tau(0)=\mathcal{R}[\rho_{0}]$ solves Problem 1, that is for all sequences $k_{0:t}$ and for all $t\geq 0$ and $\rho_{0}\in\mathfrak{D}(\mathcal{H})$
@@
for all $\rho_{0}\in\mathfrak{D}(\mathcal{H})$ and sequences $k_{0:t}$ with $t\geq 0$. By noticing that $\mathscr{N}^{\perp}\subseteq\mathscr{A}$ and using Proposition 2 we can conclude the proof. ∎
Using the same arguments as in [14], one can prove that the closure to an algebra of $\mathscr{N}^{\perp}$, orthogonal to $\mathscr{N}$ with respect to the Hilbert-Schmidt inner product ,$_{HS}$, provides the smallest algebra containing $\mathscr{N}^{\perp}$ that admits a CPTP projection, even when one allows for alternative definitions of operator product and orthogonality notions.
@@ IV-C Reduction of measurements and dynamics, separately
In the previous sections, we presented how to reduce a CE to a reduced-order one, obtaining a set of conditional dynamics that are capable of reproducing the output of interest conditioned on the outcome of the measurement. In many cases of interest, the conditional dynamic emerges as the composition of an unconditional CPTP evolution and a generalized measurement [6]. Under suitable assumptions, it is possible to obtain reduced models for both the CPTP dynamical map and the measurements, separately. In the following we thus assume that the conditional dynamics $\mathcal{M}_{k}$ are composed of a CPTP evolution map $\mathcal{E}$ following the conditioning induced by the generalized measurement $\mathcal{K}_{k}(\cdot):=M_{k}\cdot M_{k}^{\dagger}$ with $\sum_{k\in\Omega}M_{k}^{\dagger}M_{k}=\mathbb{1}$, i.e. $\mathcal{M}_{k}(\cdot)=\mathcal{E}\circ\mathcal{K}_{k}$. This ordering choice has been made to keep consistency with previous works [11, 13] but similar results can be derived if the evolution precedes the measurement.
@@
Notice that Assumption (A1) implies Assumption (A2): by definition $\mathscr{N}$ is $\mathcal{M}_{k}$-invariant for all $k$, hence for $X\in\mathscr{N}$, we have $\mathcal{E}[X]=\sum_{k}\lambda_{k}\mathcal{M}_{k}[X]\in\mathscr{N}$. This generalizes the results presented in [11, equation 3.3] and [13, Lemma 4] to the case of stochastic output matrices (not only deterministic ones). Assumption (A2) however, does not imply (A1).
@@
Under any of the assumptions (A1)-(A4), the reduced CE ({ˇ$(\{\widecheck{\mathcal{M}}_{k}\},\widecheck{\mathcal{C}})$ with conditional dynamics ˇ$\widecheck{\mathcal{M}}_{k}=\widecheck{\mathcal{E}}\widecheck{\mathcal{K}}_{k}$ where $\widecheck{\mathcal{K}}_{k}=\mathcal{R}\mathcal{K}_{k}\mathcal{J}$ and ˇ$\widecheck{\mathcal{E}}=\mathcal{R}\mathcal{E}\mathcal{J}$, $\widecheck{\mathcal{C}}=\mathcal{C}\mathcal{J}$ and $\Phi=\mathcal{R}$ solves Problem 1.
@@ V-A Measured quantum walks
Situation like this, in which conditioning induces smaller reduced models, are not possible in the classical case. In fact [11] proves that the conditional observable space always contains the unconditional one.
@@ V-B Measured Ising spin chain
The next model has been inspired by recent work on continuously-measured spin chains, see e.g. [21]. Let us consider a quantum system composed of $N\geq 4$ qubits disposed on a line [^1] We denote the operator living in $\mathfrak{B}(\mathcal{H})$ and acting as the Pauli matrix $\sigma_{j}$ with $j\in\{0,x,y,z\}$ on the $k$-th spin, with $k=1,\dots,N$ and as the identity on the rest is denoted with $\sigma_{j}^{(k)}$, i.e. $\sigma_{j}^{(k)}:=\otimes_{i=1}^{k-1}\sigma_{0}\otimes\sigma_{j}\otimes_{j=1}^{N-k}\sigma_{0}\in\mathfrak{B}(\mathcal{H})$., i.e. $\mathcal{H}\simeq(\mathbb{C}^{2})^{\otimes N}$. Let us also assume to be able to perform projective measurement of the observable $\sigma_{z}$ on the last spin, i.e. $\sigma_{z}^{(N)}$ and that, at each step, this measurement is performed with probability $1-p$ with $p\in[0,1]$ and it is left untouched with probability $p$. Under these assumptions we can model the measurement with three super operators $\mathcal{K}_{-1}(\cdot)=p\mathbb{1}\cdot\mathbb{1}$, $\mathcal{K}_{0}(\cdot)=(1-p)\Pi_{0}\cdot\Pi_{0}$ with $\Pi_{0}:=\mathbb{1}_{2^{N-1}}\otimes\left|0\middle>\middle<0\right|$ and $\mathcal{K}_{1}(\cdot)=(1-p)\Pi_{1}\cdot\Pi_{1}$ with $\Pi_{1}=\mathbb{1}_{2^{N-1}}\otimes\left|1\middle>\middle<1\right|$. Between two measurements the system evolves with a unitary evolution $\mathcal{E}(\cdot)=U\cdot U^{\dagger}$ with $U=e^{-iH}$ where $H$ is an Ising Hamiltonian, i.e.
@@
with coupling strength $\delta$. For this example, we are interested in reproducing the reduced state on the first qubit, i.e. $\tau(t)={\rm tr}_{\overline{1}}(\rho(t))$, for all initial conditions $\rho_{0}\in\mathfrak{D}(\mathcal{H})$. Here, ${\rm tr}_{\overline{1}}(\cdot)$ denotes the partial trace with respect to everything but the first qubit of the chain, which can be reconstructed from the expectation of $\left\{\sigma_{q}^{(1)},\,{q=0,x,y,z}\right\}$, representing our observables of interest. We next derive the reduced model in two cases of interest: $p=0$ and $10<p<$ 1 and we denote the spaces and maps of interest with the index 00 and $p$ to distinguish them.
@@
with $V_{k}^{\dagger}:\mathcal{H}_{k}\to\mathcal{H}$ isometries of the correct dimensions and where $\mathcal{H}_{k}=\mathcal{H}_{S,k}\otimes\mathcal{H}_{F,k}\simeq\mathbb{C}^{4}\otimes\mathbb{C}^{2^{N-3}}$. The reduced evolution map takes the form ˇ$\widecheck{\mathcal{M}}_{p}(\cdot)=\widecheck{U}_{p}\cdot\widecheck{U}_{p}^{\dagger}$ where,
$\widecheck{U}_{p}=\exp\left(-i\delta\bigoplus_{k=0}^{1}[\sigma_{0}\otimes\sigma_{z}+(-1)^{k}\sigma_{x}\otimes\sigma_{0}]\right)$
and the reduced state on the first qubit is retrieved by $\tau(t)=\sum_{k=0}^{1}{\rm tr}_{\overline{1}}[\xi_{k}(t)]$. Moreover, conditioning for the reduced model is associated to $\widecheck{\mathcal{K}_{0}}(\cdot)=p\mathbb{1}_{8}\cdot\mathbb{1}_{8}$, $\widecheck{\mathcal{K}_{0}}(\cdot)=(1-p)\hat{\Pi}_{0}\cdot\widecheck{\Pi}_{0}$ and $\widecheck{\mathcal{K}_{1}}(\cdot)=(1-p)\widecheck{\Pi}_{1}\cdot\widecheck{\Pi}_{1}$ with $\widecheck{\Pi}_{0}=\frac{\mathbb{1}_{8}+\mathbb{1}_{4}\otimes\sigma_{x}}{2}$ and $\widecheck{\Pi}_{1}=\frac{\mathbb{1}_{8}-\mathbb{1}_{4}\otimes\sigma_{x}}{2}$.
In both cases treated above, only the first two and last two qubits of a chain (of arbitrary length) influence the dynamics of the first one. Moreover, the relevant degrees of freedom in the second spin are classical, associated to the eigenbasis of $\sigma_{x}$. With this, we have that the useful information behaves like the first qubit $\mathbb{C}^{2\times 2}$ coupled classical probabilistic mixture of two qubits, i.e. $\mathbb{C}^{2\times 2}\oplus\mathbb{C}^{2\times 2}$.
VI Conclusions
We developed an algebraic method to reduce the dimension of discrete-time quantum systems, subject to repeated (generalized, imperfect) measurement and conditional dynamics, when only the expectations of certain observables have to be reproduced exactly. Possible applications of the results include: (1) obtaining more intuitive and easier-to-study models; (2) implementing more efficient simulations of quantum trajectories on both classical and quantum computers; (3) efficiently producing output samples emerging from quantum repeated-measurement models; (4) deriving reduced-state estimators to increase the capability of feedback controllers for quantum systems. Two applications models have been discussed in details, highlighting interesting features and distinguishing the quantum case from its classical analogues.
A similar approach can be developed to obtain reductions leveraging knowledge of initial conditions, extending the results of [14] to the present setting. Future developments will be devoted to extending these ideas to the continuous-time case, finding reduction bounds and symmetry-based interpretations of the obtained models, and applying the methods to concrete situations of interest.